The perimeter of a rectangle is 12 meters. Its length is twice its width. Find the dimensions of the rectangle.
Answer by checkley77(12844)

L=2W
2L+2W=12
2(2W)+2W=12
4W+2W=12
6W=12
W=12/6
W=2 ANS. FOR THE WIDTH.
L=2*2=4 ANS. FOR THE LENGTH.
PROOF:
2*4+2*2=12
8+4=12
12=12